During the year, the Senbet Discount Tire Company had gross sales of $1.12 million. The company's cost of goods sold and selling expenses were $531,000 and $221,000, respectively. The company also had debt of $860,000, which carried an interest rate of 6 percent. Depreciation was $136,000. The tax rate was 40 percent. a. What was the company's net income? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.) b. What was the company’s operating cash flow? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)
Answer #1

1)

Interest = 860,000 * 0.06 = 51,600

Net income = (Sales - COGS - selling expenses - depreciation - interest)( 1 - tax)

Net income = (1,120,000 - 531,000 - 221,000 - 136,000 - 51,600)(1 - 0.4)

Net income = $108,240

2)

We do not include interest in calculating operating cash flow but needs to include the tax saving effect of interest. Also we need to add back the depreciation.

tax expense = (1,120,000 - 531,000 - 221,000 - 136,000 - 51,600) * 0.4

tax expense = 72,160

Operating cash flow = 1,120,000 - 531,000 - 221,000 - 136,000 - 72,160 + 136,000

Operating cash flow = $295,840